OpenStudy (anonymous):

230 + 230 + 230 + 230 = A. 8120 B. 830 C. 232 D. 230 E. 226

Directrix (directrix):

230 + 230 + 230 + 230 = 4 * 230 = ?

OpenStudy (anonymous):

Uh none of your options are corresponding with the answer which is 920